OpenStudy (anonymous):

a square base pyramidal tent has a side 10ft and height 6ft. what is the volume of air inside the tent?

OpenStudy (ash2326):

Volume of Pyramid= Area of base * height Area of base=area of square height=given @meme_90 can you find the volume, you just have to plugin the values!

OpenStudy (anonymous):

is 32,000 ft^3 the answer?

OpenStudy (ash2326):

Area of square=10*10=100 ft^2 Height=6 feet Volume=100*6=600 ft^3

OpenStudy (anonymous):

\[v of \square base pyramid is b^2h/3\]

OpenStudy (ash2326):

then it'd be \[10*10*6/3=200\]

OpenStudy (anonymous):

YEP